OpenStudy (anonymous):

Help!!!! Which of the following is the slope between the points (-1, -5) and (7, -3)? A. 4 B. 1/4 C. -1/4 D. -4

OpenStudy (ranga):

Given two points (x1, y1) and (x2, y2), their slope m can be found from: m = (y2 - y1) / (x2 - x1)

OpenStudy (anonymous):

so it would be 4

OpenStudy (jdoe0001):

\(\bf \begin{array}{lllll} &x_1&y_1&x_2&y_2\\ &(-1\quad ,&5)\quad &(7\quad ,&-3) \end{array} \\\quad \\ slope = m= \cfrac{rise}{run} \implies \cfrac{y_2-y_1}{x_2-x_1}\)

OpenStudy (anonymous):

So -4

OpenStudy (jdoe0001):

hmm well.. try this one instead \(\bf slope = m= \cfrac{rise}{run} \implies \cfrac{y_2-y_1}{x_2-x_1}\implies \cfrac{(-3)-(-1)}{(7)-(-1)}\implies \cfrac{-3+1}{7+1}\)

OpenStudy (jdoe0001):

hmmm hold one... that needs some fixing

OpenStudy (jdoe0001):

\(\bf slope = m= \cfrac{rise}{run} \implies \cfrac{y_2-y_1}{x_2-x_1}\implies \cfrac{(-3)-(5)}{(7)-(-1)}\implies \cfrac{-3-5}{7+1}\)

OpenStudy (jdoe0001):

notice your "x" and "y" for each point \(\begin{array}{lllll} &x_1&y_1&x_2&y_2\\ &(-1\quad ,&5)\quad &(7\quad ,&-3) \end{array}\)

OpenStudy (anonymous):

0.25x + -4.75

OpenStudy (jdoe0001):

hmnm ohh .. .smokes... is -5.... thought it was ... 5 my bad

OpenStudy (jdoe0001):

\(\bf \begin{array}{lllll} &x_1&y_1&x_2&y_2\\ &(-1\quad ,&-5)\quad &(7\quad ,&-3) \end{array} \\\quad \\ slope = m= \cfrac{rise}{run} \implies \cfrac{y_2-y_1}{x_2-x_1}\implies \cfrac{(-3)-(-5)}{(7)-(-1)}\implies \cfrac{-3+5}{7+1}\)

OpenStudy (anonymous):

0.25

OpenStudy (anonymous):

so its 1/4

OpenStudy (jdoe0001):

yeap